The market for bread is at equilibrium at a price of 3 per loaf and a quantity of 900 loaves being sold. Given that there is no demand for bread at a price of six per loaf, and assuming that the demand curve is linear, what is the demand curve in this market?

(Please write the equation out in full in the form: Qd=a-bP, with no spaces)
1
Expert's answer
2019-09-24T08:34:55-0400

Demand curve is linear and we know its 2 points, which are (6; 0) and (3; 900).

Using this formula:

(x-x1)/(x2-x1)=(y-y1)/(y2-y1)

where (x1; y1) and (x2; y2) are points of the line,

we can find the equation for demand curve.


(x-6)/(3-6)=(y-0)/(900-0)

y=900(x-6)/(-3)

y=-300x+1800


So the answer is:

Q=-300P+1800